OpenStudy (anonymous):

Help me find the inverse? f(x) = cube root(x/8) -4

OpenStudy (jdoe0001):

\(\bf \large f(x)={\color{blue}{ y}}=\sqrt[3]{\cfrac{{\color{red}{ x}}}{8}}-4\qquad inverse\implies {\color{red}{ x}}=\sqrt[3]{\cfrac{{\color{blue}{ y}}}{8}}-4\) notice, to find the "inverse relationship" all you'd do is swap about the variables then solve for "y"

OpenStudy (nincompoop):

try obtaining the opposites - turns into + and vice versa multiplication turns into division and vice versa so roots turn into powers and vice versa see if this works

OpenStudy (anonymous):

Ok so I think the answer is f^-1(x) = 8(x + 4)^3 But the fraction in the cube root is confusing me. Is that correct?

OpenStudy (anonymous):

yeah. that's correct. god job

OpenStudy (anonymous):

Yay thank you!